# A Direct Comparison of Simultaneously Recorded Scalp, Around-Ear, and In-Ear EEG for Neural Selective Auditory Attention Decoding to Speech

### v2

@@ Introduction
EEG-based auditory attention decoding (AAD) algorithms generally fall into two categories [2]. The first category uses neural tracking, i.e., the neural signals of the listener entrain more strongly with features of the attended speech signal than with unattended ones [4, 5, 1]. These features often include acoustic ones such as the envelope or spectrogram [4, 5, 1], but also linguistic features as phoneme surprisal or semantic dissimilarity are used [6, 7]. The second category aims to directly classify attention from EEG signals, for example based on the spatial direction of attention [8, 9, 10], or using more black-box approaches based on deep learning [2, 11, 12, 13].
@@
Research towards wearable EEG systems for AAD generally take two distinct avenues: (1) a data-driven determination of optimal electrode locations based on high-density EEG, assuming ideal miniaturized EEG sensors are available, and (2) a recording system-driven approach, starting from existing wearable EEG recording systems. The data-driven approach assumes the existence of miniaturized EEG sensor nodes that can be flexibly mounted at optimally selected locations. Using data-driven channel selection and reduction techniques, Mirkovic et al. [14] (using backward elimination) and Mundanad Narayanan and Bertrand [15] (using a greedy utility-based selection) were able to reduce the number of EEG channels used for neural tracking-based algorithms from 96 to 25 and 64 to 10, respectively, without a loss in AAD performance. Furthermore, Mundanad Narayanan et al. [16] were able to reduce inter-electrode distances to , eliminating the need for long-distance montages and thereby increasing the practicality of wireless EEG sensor networks for AAD. These data-driven channel selection methods, all starting from neural tracking, typically select, as expected, electrodes above the temporal lobe, where the auditory cortex is located.
@@
However, to date, no study in the AAD literature has directly compared scalp, around-ear, and in-ear EEG. Furthermore, as noted above, in-ear AAD studies have consistently relied on an external scalp reference. To address this, we introduce a new AAD dataset in which participants attended to one of two competing talkers while their scalp, around-ear, and in-ear EEG is simultaneously recorded. This allows for a direct comparison of AAD performance across all three systems on the same data, and enables the development of tailored preprocessing pipelines and AAD algorithms for ear-based EEG systems - for example, leveraging transfer learning between scalp and ear-EEG. In this paper, we present the newly recorded AAD ear EEG-based dataset, compare the different EEG systems using a standard neural tracking algorithm, and investigate system complementarity, the influence of the reference electrode, and the potential of constructing an EEG sensor network centered around ear-based systems. To facilitate future development of ear-based AAD algorithms, the dataset is made publicly available.
@@
In this section, we provide a full description of the protocol and experimental setup for the new AAD dataset with simultaneous scalp, around-ear, and in-ear EEG recordings. The study was approved by the Institutional Review Board at Aarhus University (approval number 2024-0673174). The full dataset, along with a detailed manual and description, is publicly available [40].
@@
Each participant took part in a selective auditory attention experiment, in which they were instructed to attended to one of two simultaneously presented speech streams (left or right) while seated in a room without electric shielding. The experiment consisted of six trials, each lasting , resulting in a total of of data per participant. The two competing speech signals were presented via insert earphones, integrated into the custom earpieces used for in-ear EEG recordings (see below), and were spatially simulated at azimuth using head-related transfer functions. The signals were normalized to loudness units relative to full scale (LUFS). At the beginning of each trial, participants were presented a prelisten of the to-be-attended speech stream, followed by of competing speech signals. This prelisten fragment was intended to help participants identify the correct target stream. Figure 1 illustrates the experimental protocol. After each trial, participants answered a comprehension question about the content of the attended speech, providing motivation to remain focused. Breaks were allowed between trials at the participants’ discretion.
The speech material consisted of educational talks aimed at high school students, delivered by male Danish professors at Aarhus University. Both audio and video were available for each talk. From each talk, two non-overlapping excerpts were extracted and assigned to the set of attended and unattended stimuli, respectively. Each excerpt was used exactly and only once during the experiment. Attended excerpts were chosen such that they could be fully understood without prior context or knowledge. Full details on the talks and extracted excerpts are available online [40].
@@ Methods
We use leave-one-trial-out cross-validation (CV) for participant-specific decoding (i.e., when neural decoders are trained for a specific participant), with six folds per participant, each excluding a full trial. While such a strict CV procedure generally results in slightly lower performances than when mixing data from within a trial in both training and test set, several papers have shown the importance of a strict validation procedure in AAD to avoid overfitting and exploitation of shortcuts [11, 50]. While this is less of an issue with neural tracking-based AAD algorithms as used here, we find it important to adhere to a stricter validation procedure as a more general guideline in AAD and machine learning research. For participant-independent decoding, a leave-one-participant-out CV procedure is used.
@@
Figure 5 shows the average performance curves (mean standard error on the mean) when using a participant-independent decoder, evaluated with leave-one-participant-out CV. Using scalp EEG, the typical decrease in accuracy w.r.t. participant-specific decoding is observed, with a mean percent point decrease on windows, similarly to what was observed in Geirnaert et al. [47]. For in-ear EEG, such generalization across participants was unsuccessful, as below-significance accuracies are obtained, with a percent point decrease w.r.t. participant-specific decoding. A possible explanation is that the training set in the participant-independent case for in-ear EEG contains much more ‘noise’, contributed by those participants that did not obtain significant performance in the participant-specific case either. Moreover, differences in ear anatomy result in different personalized earpieces, with sometimes different orientations for specific EEG electrodes. Lastly, given the previous discussion on the SNR of in-ear EEG, the amount and nature of the noise could vary more between participants. Surprisingly, the around-ear EEG setup results in a much smaller decrease in performance (mean percent point) for participant-independent decoding. We hypothesize this is due to the lower inter-participant variability in electrode positioning with the standardized C-shaped electrode array, compared to the highly personalized and variable in-ear setups, while the spatial diversity is also smaller than in scalp EEG, leading to less diverse EEG patterns.
@@ Conclusion
While the current dataset reveals clear trends when comparing the different EEG configurations, the sample size of 15 participants remains relatively limited. Additionally, given the focus on the direct comparison of scalp-, around-ear, and in-ear EEG, a homogeneous group of young, normal-hearing participants was recruited. While literature has shown that linear stimulus reconstruction for AAD using scalp EEG is feasible in elderly people or individuals with hearing loss [57], future studies should consider more heterogeneous participant groups to ensure generalizability to these populations. Moreover, to assess the full potential of wearable EEG systems, the experiment could be repeated in more ecologically valid listening scenarios. In that context, the dry in-ear EEG setup used here should be compared with other dry electrode configurations. Lastly, we employed a linear stimulus reconstruction AAD algorithm. While several nonlinear (e.g., deep learning-based) AAD methods have been developed for scalp EEG, such approaches have been lacking or unsuccessful for ear-based EEG [38]. This highlights a clear need for the development of tailored nonlinear AAD algorithms for ear-based EEG.
@@
The EEG dataset generated and analyzed during the current study is publicly available in the following Zenodo repository [40], [WILL BE PUBLISHED UPON ACCEPTANCE]. In the meantime, requests for access or related questions can be directed to Simon Geirnaert (simon.geirnaert@kuleuven.be).

### v4

# A Direct Comparison of Simultaneously Recorded Scalp, Around-Ear and In-Ear EEG for Neural Selective Auditory Attention Decoding to Speech

@@ Introduction
EEG-based auditory attention decoding (AAD) algorithms generally fall into two categories [2]. The first category relies on neural tracking, i.e., the neural signals of the listener entrain more strongly with features of the attended speech signal than with unattended ones [4, 5, 1]. These features often include acoustic ones such as the envelope or spectrogram [4, 5, 1], but also linguistic features as phoneme surprisal or semantic dissimilarity are used [6, 7]. The second category aims to directly classify attention from EEG signals, for example based on the spatial direction of attention [8, 9, 10], or using more black-box approaches based on deep learning [2, 11, 12]. However, while the first category of algorithms based on neural tracking have consistently shown robustness across a wide range of datasets and setups, the second category of direct classification approaches are much more prone to overfitting, shortcut learning, and overestimated performance [2, 11, 13, 14, 15]. For this reason, we only consider the first category of algorithms based on neural tracking in this study.
@@
Research towards wearable EEG systems for AAD generally takes two distinct avenues: (1) a data-driven determination of optimal electrode locations based on high-density EEG, assuming ideal miniaturized EEG sensors are available, and (2) a recording system-driven approach, starting from existing wearable EEG recording systems. The data-driven approach assumes the existence of miniaturized EEG sensor nodes that can be flexibly mounted at optimally selected locations. Using data-driven channel selection and reduction techniques, Mirkovic et al. [16] (using backward elimination) and Mundanad Narayanan and Bertrand [17] (using a greedy utility-based selection) were able to reduce the number of EEG channels used for neural tracking-based algorithms from 96 to 25 and 64 to 10, respectively, without a loss in AAD performance. Furthermore, Mundanad Narayanan et al. [18] were able to reduce inter-electrode distances to , eliminating the need for long-distance montages and thereby increasing the practicality of wireless EEG sensor networks for AAD. These data-driven channel selection methods, all starting from neural tracking, typically select, as expected, electrodes above the temporal lobe, where the auditory cortex is located.
@@
However, to date, no study in the AAD literature has directly compared scalp, around-ear, and in-ear EEG. Furthermore, as noted above, in-ear AAD studies have consistently relied on an external scalp reference. To address this, we introduce a new AAD dataset in which participants attended to one of two competing talkers while their scalp, around-ear, and in-ear EEG is simultaneously recorded. This allows for a direct comparison of AAD performance across all three systems on the same data, and enables the development of tailored preprocessing pipelines and AAD algorithms for ear-based EEG systems - for example, leveraging transfer learning between scalp and ear-EEG. In this paper, we present the newly recorded AAD ear EEG-based dataset, compare the different EEG systems using a standard neural tracking algorithm, and investigate system complementarity, the influence of the reference electrode, and the potential of constructing an EEG sensor network centered around ear-based systems. To facilitate future development of ear-based AAD algorithms, the dataset is made publicly available [42].
@@
In this section, we provide a full description of the protocol and experimental setup for the new AAD dataset with simultaneous scalp, around-ear, and in-ear EEG recordings. The study was approved by the Institutional Review Board at Aarhus University (approval number 2024-0673174) and conducted following the applicable guidelines and regulations. The full dataset, along with a detailed manual and description, is publicly available [42].
@@
Each participant took part in a selective auditory attention experiment, in which they were instructed to attended to one of two simultaneously presented speech streams (left or right) while seated in a room without electric shielding. The experiment consisted of six trials, each lasting , resulting in a total of of data per participant. The two competing speech signals were presented via insert earphones, integrated into the custom earpieces used for in-ear EEG recordings (see below), and were spatially simulated at azimuth using head-related transfer functions. The signals were normalized to loudness units relative to full scale (LUFS). At the beginning of each trial, participants were presented a prelisten of the to-be-attended speech stream, followed by of competing speech signals. This prelisten fragment was intended to help participants identify the correct target stream. Figure 1 illustrates the experimental protocol. After each trial, participants answered an open-ended comprehension question requiring a short, precise answer about the content of the attended speech. Performance on these questions was not logged, nor used as an exclusion criterion, but was only included to help maintain the participant’s attention and motivation. Breaks were allowed between trials at the participants’ discretion.
The speech material consisted of six educational talks aimed at high school students, delivered by six different male Danish professors at Aarhus University. Both audio and video were available for each talk. From each talk, two non-overlapping excerpts were extracted and assigned to the set of attended and unattended stimuli, respectively, with each excerpt used exactly and only once during the experiment. In total, there were thus six different speakers to attend to, reducing potential bias due to familiarization with a specific speaker or voice. Attended excerpts were chosen such that they could be fully understood without prior context or knowledge. Unlike many other AAD studies that truncate silences to a maximal duration, the speech signals here were not further manipulated, i.e., silences were left at their natural length. Across all stimuli, silences longer than had a mean ( standard deviation) duration of . As silence duration is known to affect onset processing and neural tracking correlations of envelopes [43], this difference should be considered when comparing with other AAD studies. Full details on the talks and extracted excerpts are available online [42].
@@ Methods
We use leave-one-trial-out cross-validation (CV) for participant-specific decoding (i.e., when neural decoders are trained for a specific participant), with six folds per participant, each excluding a full trial. While such a strict CV procedure generally results in slightly lower performances than when mixing data from within a trial in both training and test set, several papers have shown the importance of a strict validation procedure in AAD to avoid overfitting and exploitation of shortcuts [11, 13, 14]. While this is less of an issue with neural tracking-based AAD algorithms as used here, we find it important to adhere to a stricter validation procedure as a more general guideline in AAD and machine learning research. Furthermore, the leave-one-trial-out CV guarantees that no information leakage from the training to the test set is possible due to the per-trial preprocessing. For participant-independent decoding, a leave-one-participant-out CV procedure is used.
@@
To assess the impact of EEG preprocessing, especially for ear-based setups, we re-evaluate AAD accuracies on decision windows (Figure 3(b)) after omitting most artifact removal steps, namely EOG regression (for scalp EEG), bad channel removal, ASR, and high-power artifact rejection. The average AAD accuracy increases to for scalp EEG (all 15 participants significant, with and without Benjamini–Hochberg correction), but decreases to for around-ear EEG (11 participants significant with and without correction) and to for in-ear EEG (with 4/7 significant accuracies with/without correction). The substantial performance drop in AAD accuracy and number of participants with significant performance for both around-ear and in-ear EEG demonstrates the higher importance of more rigorous artifact removal for achieving reliable decoding with ear-based EEG. All results reported in the remainder of this paper are therefore based on the full preprocessing pipeline.
@@
Figure 5 shows the average performance curves (mean standard error on the mean) when using a participant-independent decoder, evaluated with leave-one-participant-out CV. Using scalp EEG, the typical decrease in accuracy w.r.t. participant-specific decoding is observed, with a mean percent point decrease on windows, similarly to what was observed in Geirnaert et al. [50]. For in-ear EEG, such generalization across participants was unsuccessful, as below-significance accuracies are obtained, with a percent point decrease w.r.t. participant-specific decoding. A possible explanation is that the training set in the participant-independent case for in-ear EEG contains much more ‘noise’, contributed by those participants that did not obtain significant performance in the participant-specific case either. For example, using only the six participants that show significant AAD accuracy with in-ear EEG (after correction), a AAD accuracy on windows could be achieved with in-ear EEG (3 out of 6 participants significant). Moreover, differences in ear anatomy result in different personalized earpieces, with sometimes different orientations for specific EEG electrodes. Lastly, given the previous discussion on the SNR of in-ear EEG, the amount and nature of the noise could vary more between participants. Surprisingly, the around-ear EEG setup results in a much smaller decrease in performance (mean percent point) for participant-independent decoding. We hypothesize this is due to the lower inter-participant variability in electrode positioning with the standardized C-shaped electrode array, compared to the highly personalized and variable in-ear setups, while the spatial diversity is also smaller than in scalp EEG, leading to less diverse EEG patterns.
@@ Conclusion
While the current dataset reveals clear trends when comparing the different EEG configurations, the sample size of 15 participants remains relatively limited. Additionally, given the focus on the direct comparison of scalp-, around-ear, and in-ear EEG, a homogeneous group of young, normal-hearing participants was recruited. While literature has shown that linear stimulus reconstruction for AAD using scalp EEG is feasible in elderly people or individuals with hearing loss [59], future studies should consider more heterogeneous participant groups to ensure generalizability to these populations. Moreover, to assess the full potential of wearable EEG systems, the experiment could be repeated in more ecologically valid listening scenarios. In that context, the dry in-ear EEG setup used here should be compared with other dry electrode configurations. Another limitation is that only one comprehension question was asked after each -trial. This might not sufficiently motivate participants to sustain focus throughout the trial. Since task performance was probed only at the end of each trial and responses were not logged, fluctuations in participants’ attention could not be accounted for. Lastly, we employed a linear stimulus reconstruction AAD algorithm. While several nonlinear (e.g., deep learning-based) AAD methods have been developed for scalp EEG, such approaches have been lacking or unsuccessful for ear-based EEG [40]. This highlights a clear need for the development of tailored nonlinear AAD algorithms for ear-based EEG.
@@
The EEG dataset generated and analyzed during the current study is publicly available in the following Zenodo repository [42]: https://doi.org/10.5281/zenodo.16536441.